### v1

@@ 3.4.2 Test of Proposition 4: Pass-Through and Bilateral Market Shares
Table 3 presents the results using the baseline fixed effects. Column (1) shows that, on average, pass-through into duty-exclusive prices is incomplete: a 10% tariff increase reduces exporter prices by 1.5% , corresponding to 85% pass-through rate. Column (2) adds an interaction with relationship age to account for the role of match longevity, which has been shown to influence price adjustments. The results confirm that pass-through rises with relationship length, consistent with the results in Heise (2024).
@@
Columns (3)–(6) show that the coefficient on supplier share ($\alpha_{s}$) is positive but insignificant, suggesting limited pass-through heterogeneity on the exporter side. In contrast, the coefficient on buyer share ($\alpha_{x}$) is consistently negative and statistically significant, indicating that importers with greater buyer share face lower pass-through. Table D.2 in Appendix D.2 confirms that these patterns hold under more demanding fixed effects. Table D.3 shows they are robust to alternative price definitions and general equilibrium controls.
@@
To assess nonlinearities, we interact tariff changes with quartiles of lagged supplier and buyer shares. Figure D.1 shows no systematic pattern across supplier share quartiles but a clear, monotonic decline across buyer share quartiles, regardless of fixed effects.
Together, these results suggest that pass-through is largely unresponsive to supplier concentration but declines strongly with buyer concentration. This pattern aligns with the model’s predictions under two-sided market power: the strong buyer share gradient provides direct evidence of decreasing returns, as formalized in Proposition 4, while the weak supplier share gradient reflects the dominance of cost-channel, which suggests a high-$\phi$, low-$\theta$ environment.
4 Structural Estimation
The patterns documented above are consistent with the model’s predictions under two-sided market power. However, as discussed, they do not warrant direct inference on the structural parameters. We now turn to a structural estimation approach to quantify the role of two-sided market power in shaping international prices.
Our estimation targets two key parameters: the importer’s bargaining power, $\phi$, and the returns to scale parameter, $\theta$, which governs the elasticity of foreign export supply. The remaining parameters are taken from the literature or directly measured from the data.
We set the elasticity of substitution across foreign varieties to $\rho=10$, consistent with Anderson and van Wincoop (2004) and Edmond et al. (2023), who adopt similar values to match observed U.S. markups. The downstream demand elasticity faced by importers is set to $\nu=4$, based on the estimates in Broda and Weinstein (2006). [^23] Appendix E.1 provides further discussion. The elasticity of importer $j$’s marginal cost with respect to the foreign input price index is set to $\gamma=0.5$, following Eldridge and Powers (2018), who document the share of imported inputs in total material costs for U.S. manufacturers. Input cost shares $\alpha_{jht}$ are directly measured from the data.
Finally, we set the importer’s returns to scale parameter to $\varrho=1$ in the baseline specification. As discussed in Section 2.3.1, the parameters $\nu$, $\gamma$, and $\varrho$ jointly determine the elasticity $\eta$ of importer $j$’s foreign input bundle $q^{f}_{j}$ with respect to its price index $p^{f}_{j}$. Since $\varrho$ enters the model only through $\eta$, fixing it to one does not restrict generality. We verify that our results are robust to alternative values of $\eta$.
@@ 4.1 Identification and Estimation of the Parameters $\theta$ and $\phi$
Let $\Omega_{ijt}$ denote the information set available to a given $i-j$ pair during negotiations. This includes observed market shares ($s_{ijht}$, $x_{ijht}$, $\alpha_{jht}$) and calibrated parameters ($\nu$, $\gamma$, $\rho$, $\varrho$). As shown in equation (2.6), the bilateral markup depends only on the model primitives $(\phi,\theta)$ conditional on $\Omega_{ijt}$, i.e., $\mu_{ij}=\mu\left(\phi,\theta\mid\Omega_{ijt}\right).$ The log price of product $h$ transacted between exporter $i$ and importer $j$ in year $t$ can thus be written as:
@@
$\ln p_{ijht}=\ln\mu\left(\phi,\theta\mid\Omega_{ijt}\right)+\ln c_{iht},$
where $\ln c_{iht}$ denotes the exporter’s marginal cost. From equation (2.1), this is given by:
$\ln c_{ijht}=\frac{1-\theta}{\theta}\ln q_{iht}+\ln k_{iht}.$
In the theoretical model, the term $k_{iht}$ captured exporter-level cost shifters. In the empirical implementation, we generalize this term to allow for match-specific cost components, such as relationship-specific know-how or specialization, by letting it vary flexibly at the match-year level, $k_{ijht}$, thereby absorbing all $(i,j,h,t)$-specific variation. In contrast, the term $\frac{1-\theta}{\theta}\ln q_{iht}$ is constant across importer matches for a given supplier–product–year. Identification of $(\phi,\theta)$ thus relies on cross-sectional variation in bilateral market shares, as discussed next.
Consider an exporter $i$ matched with two importers, $j$ and $\ell$. Conditional on the joint information set $\boldsymbol{\Omega}_{ij\ell t}\equiv(\Omega_{ijt},\Omega_{i\ell t})$, we assume that the unobserved component of marginal cost is mean-independent of the buyer identity: $\Delta k_{ij\ell ht}\equiv\mathbb{E}_{k}\left[k_{ijht}-k_{i\ell ht}\mid\boldsymbol{\Omega}_{ij\ell t}\right]=0.$ [^24] Omitted variables may induce $\Delta k_{ij\ell ht}\neq 0$, raising endogeneity concerns. One possible issue is endogenous network formation: unobserved factors may simultaneously affect both match formation and pricing, biasing estimates of $\hat{\phi}$ and $\hat{\theta}$. We address this by differencing across buyers of the same exporter, which removes exporter-level shocks common to all matches. To address remaining concerns, we further implement an instrumental variable strategy. Taking log price differences across buyers $j$ and $\ell$ served by the same exporter $i$ in year $t$ yields the following moment condition:
$g(\phi,\theta\mid\boldsymbol{\Omega}_{ij\ell t})\equiv\mathbb{E}_{k}\Big{[}\ln p_{ijht}-\ln p_{i\ell ht}-\big{(}\ln\mu(\phi,\theta\mid\Omega_{ijt})-\ln\mu(\phi,\theta\mid\Omega_{i\ell t})\big{)}\Big{|}\boldsymbol{\Omega}_{ij\ell t}\Big{]}=0.$ (4.1)
Identification requires that equation (4.1) does not hold for two pairs ($\phi,\theta)$ such that $(\phi^{A},\theta^{A})\neq(\phi^{B},\theta^{B})$. Since the oligopoly markup is independent of the returns to scale parameter $\theta$, identification of $\theta$ requires that the oligopsony channel plays a role in price determination, i.e., $\phi>0$, a condition supported by the reduced-form evidence in Tables 2 and 3.
@@
We therefore focus on the empirically relevant case of bilateral bargaining power $\phi\in(0,1)$. The markup function $\mu(\phi,\theta\mid\Omega)$ is strictly monotonic in both parameters and thus invertible in each. It follows that the moment condition in equation (4.1) is also invertible in $\phi$ and $\theta$. Identification then relies on observing multiple importer–exporter pairs in the same year, or multiple matches for a given exporter over time, under the assumption that bargaining weights remain constant across matches. [^25] Formally, identification relies on the nonlinearity of the markup equation (2.6) in $s_{ijht}$ and $x_{ijht}$. Consider moment conditions from two periods $t$ and $t-1$: the associated derivatives with respect to $(\phi,\theta)$ are not collinear, satisfying the full-rank condition. Similar variation across multiple matches for the same exporter in a given year (e.g., $i$$j$$k$ vs. $i$$j$$\ell$) also secures identification.
@@
Importantly, our identification strategy does not hinge on strong assumptions about the exogeneity of tariffs or other aggregate variables. While the estimated parameters may be shaped by general equilibrium forces, identification relies solely on cross-sectional variation across buyers within supplier–product–year cells. As a result, we do not need to specify or model the broader general equilibrium environment.
@@
where $\mathbf{g}(\mathbf{\phi},\theta)$ stacks all moment conditions in equation (4.1) across all $i$$j$$\ell$ pairs and years and $\mathbf{W}$ is the optimal weighting matrix. [^26] Appendix E.2 presents Monte Carlo simulations based on a data-generating process that mirrors the setup in Section 2. The results confirm that our estimators is consistent.
@@
In particular, the vector $\mathbf{Z}$ includes the total number of importers and exporters in each HS10 product-year, which we interpret as proxies for the pool of potential US buyers and foreign suppliers in a given variety. We also include in $\mathbf{Z}$ the mean and median of the distributions of the two bilateral shares within each year, excluding the focal pairs $i-j$ and $i-\ell$ to preserve over-identification. These instruments vary with the competitive structure within each HS10 product-year and are correlated with the endogenous variables through market structure, but, by construction, are not correlated with the idiosyncratic shocks affecting individual matches.
@@
While our baseline assumes a constant bargaining weight $\phi$ across all importer–exporter pairs, we also consider an extension to allow $\phi$ to vary at the pair level.
@@
where $\boldsymbol{\kappa}$ is a parameter vector to be estimated and $\mathbf{X}_{ijht}$ includes covariates that plausibly influence bargaining outcomes but are not direct determinants of gains from trade in our model. Specifically, we include: (i) the longevity of the $i-j$ relationship, (ii) the number of transactions between $i-j$ in a year, (iii) the relative outside option of the two, measured by the ratio of the quantity of the exporter $i$’s sales to buyers other than $j$ in year $t-1$ over the quantity of the importer $j$’ purchases from suppliers other than $i$ in year $t-1$, and (iv) an indicator variable of whether the buyer and supplier transact multiple HS10 products.
@@ 4.2 Estimation Results
We estimate equation (4.2) using data from 2001 to 2016. We exclude 2017 and 2018, as these years will be used to validate the model out-of-sample in Section 4.3, leveraging the tariff shocks that occurred during this period. To avoid convergence issues when $\phi$ is near one, we estimate the transformed parameter $\bar{\phi}\equiv\ln\frac{\phi}{1-\phi}$, which enters the markup equation linearly.
@@
Panel A: Calibrated Parameters $\hat{\nu}$ $\hat{\gamma}$ $\hat{\rho}$ 4 0.5 10 Panel B: Estimated Parameters (GMM) (1) (2) (3) (4) Rel. bargaining power: $\ln\widehat{\frac{\phi}{1-\phi}}$ 1.565 0.863 (0.055) (0.043) Returns to scale ($\hat{\theta}$) 0.454 0.497 0.383 0.502 (0.004) (0.006) (0.006) (0.007) Constant 4.118 1.454 (0.428) (0.180) Longevity -0.360 0.332 (0.062) (0.064) Number of HS10 transactions -0.264 -0.003 (0.029) (0.014) Multiple HS10 dummy -0.180 0.131 (0.047) (0.034) Lagged outside option -0.235 -0.230 (0.031) (0.030) None Yes Yes No No $\text{FE}_{h}+\text{FE}_{t}+\text{FE}_{j}$ No No Yes Yes Observations 3,120,000 Panel C: Implied Bargaining Powers ($\hat{\phi}$) Mean 0.827 0.922 0.703 0.860 (0.008) (0.074) (0.009) (0.099) Median 0.945 0.886 (0.074) (0.099) Notes: This table presents model estimates based on our main estimation sample, which focuses on U.S. imports of intermediate inputs and capital goods for the period 2001-2016. Panel A reports calibrated parameters: the elasticity of demand ($\nu$), the elasticity of costs with respect to foreign input prices ($\gamma$), and the elasticity of substitution across foreign varieties ($\rho$). We set $\varrho=1$, so that $\eta=2.5$. Panel B presents GMM estimates. Columns (1) and (3) impose a constant $\phi$ across bilateral pairs, while Columns (2) and (4) estimate the full vector $\boldsymbol{\kappa}$ to allow for heterogeneity in bargaining power. Specifications differ in the inclusion of fixed effects. Controls include: (i) the log of relationship longevity between exporter $i$ and importer $j$; (ii) the log of the number of transactions between $i$ and $j$ in a given year; (iii) the log of the relative outside option, defined as the ratio of exporter $i$’s sales to other U.S. buyers (excluding $j$) over importer $j$’s purchases from other suppliers (excluding $i$), both in year $t{-}1$; and (iv) a dummy variable equal to one if the $i-j$ pair transacts in more than one HS10 product. Panel C reports the mean and median of the implied bargaining power. Standard errors are robust; those in Panel C are computed using the delta method. The set of instruments includes the number of exporters and importers at the HS10 level, as well as lagged bilateral shares (excluding the focal pair). The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
@@
The parameters are precisely estimated. Across specifications, U.S. importers appear to wield substantial bargaining power, with estimated values of $\phi$ ranging from 0.70 to 0.92. Our preferred estimate, reported in Column (1), is $\hat{\phi}=0.83$, implying that U.S. importers have, on average, roughly four times the bargaining power of their foreign suppliers. [^27] These findings are consistent with evidence from related settings. Morlacco (2019) documents significant markdowns by French importers in input trade, while Atkin et al. (2024) show that Argentine importers often exercise considerable bargaining power.
The returns to scale parameter $\hat{\theta}$ is consistently estimated below one, ranging from 0.40 to 0.50 across specifications, with a preferred estimate of 0.45.282828While standard estimates of returns to scale often cluster near one, $\theta$ here captures a short-run returns to scale elasticity, i.e., the slope of the exporter’s marginal cost curve during bargaining. This implies a residual export supply elasticity between 0.25 and 0.375 for the average importer, indicating relatively steep foreign supply curves. [^29] The implied residual supply elasticity is computed from the marginal cost slope $c^{\prime}_{q_{ijh}}=\frac{1-\theta}{\theta}\cdot x_{ijh}$, using the average buyer share of 0.25 from Table 1. These values are consistent with evidence from U.S. manufacturing under short-run constraints: Boehm and Pandalai-Nayar (2022) report median inverse elasticities around 0.3 at typical capacity levels, and Broda et al. (2008) document similarly low elasticities across many traded goods.
Moving to the estimates of the vector $\hat{{\boldsymbol{\kappa}}}$, we find that the coefficients on relationship longevity, frequency of transactions, and the multiple-product indicator are highly significant, although their sign varies depending on the set of fixed effects included. By contrast, the coefficient on the relative outside option is stable across specifications: importers hold less bargaining power when their supplier has a stronger outside option. Specifically, an increase in the supplier’s past sales to other buyers relative to the importer’s purchases from other suppliers is consistently associated with lower bargaining power for the importer.
@@
We assess the robustness of our structural estimates to alternative sample definitions and model calibrations. Appendix D.3 first considers a broader sample that includes all products in the BEC classification, notably extending the baseline by adding consumption goods. We also examine an alternative sample that uses related-party trade indicators from the LFTTD instead of ORBIS. On the calibration side, we consider a lower elasticity of substitution across foreign varieties ($\rho=5$ instead of 10) and introduce decreasing returns to scale in downstream production by setting $\varrho=0.5$ rather than 1. This choice aligns with the estimated returns to scale on the exporter side and allows us to test the robustness of the estimates to alternative values of the importer’s downstream demand elasticity, $\eta$. Across all variations, the structural estimates remain highly stable.
@@
Using equation (2.6), we compute markups for all buyer–supplier–product matches given the estimated parameters and the observed distribution of market shares.
The resulting markup distribution is clustered near the competitive benchmark. Our preferred estimates yield a mean markup of 0.94, with the median even closer to competitive levels. These low markups reflect the strong countervailing power of importers. The model implies that observed markups are a convex combination of oligopoly and oligopsony markups, which average 1.34 and 0.87, respectively. The large estimated bargaining power of buyers ($\hat{\phi}\approx 0.8$) shifts weight toward the oligopsony case, allowing importers to extract a substantial share of the surplus from exporters. [^30] The effective bargaining weight $\omega_{ijh}$ averages 0.77 (standard deviation 0.05), slightly below $\hat{\phi}$. This indicates that network effects, on average, dampen the importers’ effective bargaining power relative to $\phi$, although the gap is small.

### v2

@@ 3.4 Reduced-Form Evidence on Prices and Pass-Through
Table 3 presents the results using the baseline fixed effects. Column (1) shows that, on average, pass-through into duty-exclusive prices is incomplete: a 10% tariff increase reduces exporter prices by 1.3% , corresponding to 87% pass-through rate. Column (2) shows that this finding is robust to controlling for relationship age. It also reveals that pass-through increases with the length of the relationship, consistent with heise2024firm.
@@
Columns (3), (5), and (6) show that the coefficient on supplier share ($\alpha_{s}$) is positive but statistically insignificant, suggesting limited heterogeneity in pass-through on the exporter side. In contrast, the coefficient on buyer share ($\alpha_{x}$) (Columns (4)-(6)) is negative and highly statistically significant, indicating that importers with greater buyer share experience lower pass-through. Table D.2 in the Appendix D.2 confirms that these patterns persist under more demanding fixed effects, while Table D.3 shows they are robust to alternative price definitions and the inclusion of general equilibrium controls.
@@
To assess for nonlinearities, we interact tariff changes with quartiles of supplier and buyer shares. Figure D.1 shows no systematic pattern across supplier share quartiles, but a clear, monotonic decline across buyer share quartiles, regardless of the fixed effects.
Taken together, these results indicate that pass-through is largely insensitive to supplier concentration but declines sharply with buyer concentration. By Proposition 4, the negative gradient in $x$ implies $\theta<1$; these patterns are consistent with the first row of Figure 2.
4 Structural Evidence
The reduced-form patterns in Section 3.4 are consistent with two-sided market power and decreasing returns, but they cannot separately identify the structural parameters or convincingly isolate the model’s mechanisms from general-equilibrium effects. We therefore adopt a structural approach. Estimation focuses on $\phi$ and $\theta$, the two primitives that govern bargaining and decreasing returns; the remaining parameters are calibrated from the literature.
We set the elasticity of substitution across foreign varieties to $\rho=10$, consistent with Anderson2004 and edmond2023costly, who use similar values to match observed U.S. markups. The downstream demand elasticity is set to $\nu=4$, based on variety-level import demand estimates in Broda2006.232323Appendix E.1 provides additional details.
The baseline model features a single foreign input, whereas firms in the data import many (Section 3.2). In the empirical implementation, we therefore replace the common elasticity $\gamma$ with the term $\alpha_{jh}\gamma$, where $\alpha_{jh}$ is the observed expenditure share of input $h$ in firm $j$’s imported intermediates. We set $\gamma=0.5$ equal to the share of imported inputs in total material costs for U.S. manufacturers from Eldridge2018.
Finally, the importer’s returns-to-scale parameter $\varrho$ has no direct counterpart in the supply-chain production literature. We therefore adopt $\varrho=1$ as a parsimonious benchmark and assess robustness to alternative values, including $\varrho=0.5$. Because equilibrium outcomes depend on importer technology $(\nu,\gamma,\varrho)$ only through the composite elasticity $\eta_{jh}$, these exercises are equivalent to varying $\eta_{jh}$ directly.
@@ 4.1 Identification and Estimation of the Parameters $\theta$ and $\phi$
Let $\Omega_{ijt}$ denote the information set available to pair $(i,j)$ at the time of negotiation. This includes observed market shares and calibrated parameters. As shown in equation (2.10), conditional on $\Omega_{ijt}$ the bilateral markup depends only on the structural parameters $(\phi,\theta)$, so that $\mu_{ij}=\mu\!\left(\phi,\theta\mid\Omega_{ijt}\right).$
We write the log price of input $h$ sold by exporter $i$ to importer $j$ in year $t$ as
@@
$\ln p_{ijht}=\ln\mu\!\left(\phi,\theta\mid\Omega_{ijt}\right)+\ln MC_{ijht}\qquad\text{with}\qquad\ln MC_{ijht}=\frac{1-\theta}{\theta}\,\ln q_{iht}+\ln k_{ijht},$
where $p_{ijht}$ is the CIF duty-exclusive unit value and $k_{ijht}$ denotes a match-year-specific cost component.
In the theoretical model, the marginal cost term ($\ln MC_{ijht}$) is common across buyers of a given exporter-product-year. In the empirical implementation, we allow it to vary flexibly at the match-year level to absorb any $(i,j,h,t)$-specific cost differences in levels. The scale term $\frac{1-\theta}{\theta}\ln q_{iht}$ remains common across buyers within a supplier-product-year and captures the effect of decreasing returns.
Consider exporter $i$ matched with two importers, $j$ and $\ell$. Conditional on the joint information set $\boldsymbol{\Omega}_{ij\ell t}\equiv(\Omega_{ijt},\Omega_{i\ell t})$, we assume mean independence of the unobserved cost component across buyers:
$\mathbb{E}_{k}\!\left[\ln k_{ijht}-\ln k_{i\ell ht}\mid\boldsymbol{\Omega}_{ij\ell t}\right]=0.$ (4.1)
Differencing log prices across buyers served by the same exporter yields the moment condition
$\mathbb{E}_{k}\Big[\ln p_{ijht}-\ln p_{i\ell ht}-\big(\ln\mu(\phi,\theta\mid\Omega_{ijt})-\ln\mu(\phi,\theta\mid\Omega_{i\ell t})\big)\mid\boldsymbol{\Omega}_{ij\ell t}\Big]=0.$ (4.2)
After differencing, exporter-level marginal cost components cancel out, so the remaining price differences across buyers reflect differences in bilateral markups. Since the markup function depends only on $(\phi,\theta)$ conditional on $\Omega_{ijt}$, variation in market shares identifies the structural parameters.
Identification requires that equation (4.2) holds uniquely at the true $(\phi,\theta)$. Because the oligopoly component of the markup does not depend on $\theta$, identification of $\theta$ requires $\phi>0$, so that the oligopsony channel affects pricing. The empirical evidence in Tables 2 and 3 supports the presence of such bilateral effects.
@@
For $\phi\in(0,1)$, the markup function is strictly monotonic in both parameters over the empirically relevant range. This monotonicity implies local invertibility of the moment condition in $(\phi,\theta)$. Identification is therefore secured by observing multiple matches for the same exporter, either within a year or over time, under the assumption that bargaining weights are constant across matches. [^24] Identification follows from the nonlinearity of equation (2.10) in $s_{ijht}$ and $x_{ijht}$, which ensures that the derivatives with respect to $(\phi,\theta)$ satisfy a rank condition.
Threats to Identification
A first concern is endogenous match formation or omitted match characteristics. If such factors affect $\Delta k_{ij\ell ht}$, assumption (4.1) may be violated, potentially biasing estimates of $\hat{\phi}$ and $\hat{\theta}$. Our strategy mitigates this concern by differencing across buyers of the same exporter, which removes shocks common to all matches of that supplier. To further address residual concerns, we implement an instrumental variable strategy; see Appendix E.2 for details.
A second concern is that price dispersion within supplier-product-year cells, even within narrowly defined HS10 categories, may reflect product or service differences rather than bargaining. In the model, such heterogeneity is captured by the idiosyncratic cost component $k_{ijht}$, which shifts price levels but does not drive identification. Identification instead comes from the systematic relationship between prices and bilateral market shares. To overturn our interpretation, unobserved match characteristics would need to replicate the nonlinear price-share relationship generated by bilateral bargaining under decreasing returns.
@@
Finally, identification does not rely on strong assumptions about the exogeneity of tariffs or other aggregate shocks. Although general equilibrium forces may influence the estimated parameters, identification is based solely on cross-sectional variation across buyers within supplier-product-year cells. We therefore require only that such aggregate forces do not differentially affect prices across matches of the same supplier.
@@
[…]
@@
In particular, the vector $\mathbf{Z}$ includes the total number of importers and exporters in each HS10 product-year, which we interpret as proxies for the pool of potential US buyers and foreign suppliers in a given variety. We also include in $\mathbf{Z}$ the mean and median of the distributions of the two bilateral shares within each year, excluding the focal pairs iji-j and $i$-$\ell$ to preserve over-identification. These instruments vary with the competitive structure within each HS10 product-year and are correlated with the endogenous variables through market structure, but, by construction, are not correlated with the idiosyncratic shocks affecting individual matches.
@@
While our baseline assumes a constant bargaining weight $\phi$ across all importer-exporter pairs, we also consider an extension to allow $\phi$ to vary at the pair level.
@@
where $\boldsymbol{\kappa}$ is a parameter vector to be estimated and $\mathbf{X}_{ijht}$ includes covariates that plausibly influence bargaining outcomes but are not direct determinants of gains from trade in our model. Specifically, we include: (i) the longevity of the iji-j relationship, (ii) the number of transactions between iji-j in a year, (iii) the supplier’s outside option relative to the importer’s, measured by the ratio of the quantity of the exporter $i$’s sales to buyers other than $j$ in year t1t-1 over the quantity of the importer $j$’ purchases from suppliers other than $i$ in year t1t-1, and (iv) an indicator variable of whether the buyer and supplier transact multiple HS10 products.
@@ 4.2 Estimation Results
We estimate equation (4.3) using data from 2001 to 2016. We exclude 2017 and 2018, which we reserve for out-of-sample validation in Section 4.3 using the tariff shocks during that period. To avoid convergence issues when $\phi$ is close to one, we estimate the transformed parameter $\bar{\phi}\equiv\ln\frac{\phi}{1-\phi}$, which enters the markup equation linearly.
@@
Panel A: Calibrated Parameters $\hat{\nu}$ $\hat{\gamma}$ $\hat{\rho}$ 4 0.5 10 Panel B: Estimated Parameters (GMM) (1) (2) (3) (4) Rel. bargaining power: $\ln\widehat{\frac{\phi}{1-\phi}}$ 1.563 1.168 (0.051) (0.043) Returns to scale ($\hat{\theta}$) 0.455 0.505 0.417 0.508 (0.004) (0.004) (0.005) (0.006) Constant 4.327 2.099 (0.361) (0.231) Longevity -0.216 0.476 (0.043) (0.090) Number of HS10 transactions -0.342 -0.115 (0.030) (0.024) Multiple HS10 dummy 0.004 0.113 (0.038) (0.039) Lagged outside option -0.285 -0.242 (0.028) (0.032) None Yes Yes No No $\text{FE}_{h}+\text{FE}_{t}+\text{FE}_{j}$ No No Yes Yes Observations 3,120,000 Panel C: Implied Bargaining Powers ($\hat{\phi}$) Mean 0.827 0.931 0.763 0.904 (0.007) (0.081) (0.008) (0.079) Median 0.957 0.927 (0.081) (0.079) Notes: This table presents model estimates using our main estimation sample of U.S. imports of intermediate inputs and capital goods, 2001–2016. Panel A reports calibrated parameters: demand elasticity ($\nu$), cost elasticity to foreign input prices ($\gamma$), and the elasticity of substitution across foreign varieties ($\rho$). Panel B reports GMM estimates. Columns (1) and (3) impose a constant $\phi$, while Columns (2) and (4) allow heterogeneous bargaining power via $\boldsymbol{\kappa}$. Specifications differ in fixed effects. Controls include relationship longevity, transaction intensity, the relative outside option (lagged), and a multi-product indicator. Panel C reports the mean and median implied bargaining power. Robust standard errors; Panel C uses the delta method. Instruments include HS10-level counts of exporters and importers and lagged bilateral shares (excluding the focal pair). All estimations use CIF duty-exclusive unit values. The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
@@
The parameters are precisely estimated. Across specifications, U.S. importers appear to wield substantial bargaining power, with estimated values of $\phi$ ranging from 0.7 to 0.9. Our preferred estimate, reported in Column (1), is $\hat{\phi}=0.83$, implying that U.S. importers have, on average, roughly four times the bargaining power of their foreign suppliers.
The returns to scale parameter $\hat{\theta}$ is consistently estimated below one, ranging from 0.42 to 0.51 across specifications, with a preferred estimate of 0.46.252525While our estimate of $\theta$ lies below standard production-function estimates, which are typically closer to one, this reflects the short-run nature of the object we identify. Standard estimates rely on firm-level variation in inputs and outputs and need not recover the same object as our parameter, which is a strictly short-run parameter. This implies an inverse residual export supply elasticity of about 0.24-0.35 for the average importer. [^26] The implied elasticity is computed as $d\ln MC_{ih}/d\ln q_{ijh}=\frac{1-\theta}{\theta}\cdot x_{ijh}$, using the average buyer share of 0.25 from Table 1. These values are consistent with evidence from U.S. manufacturing under short-run constraints: boehm2022convex report median inverse elasticities around 0.3 at typical capacity levels, and broda2008optimal document similarly low elasticities across many traded goods.
Moving to the estimates of the vector $\hat{{\boldsymbol{\kappa}}}$, the results are mixed across the relationship-level controls. Relationship longevity is significant in both specifications but changes sign when fixed effects are included, suggesting that its correlation with bargaining power is sensitive to how unobserved heterogeneity is handled. The coefficient on transaction frequency is consistently negative and significant, while the multiple-product indicator is positive but precisely estimated only in the fixed-effects specification. By contrast, the coefficient on the relative outside option is stable and precisely estimated across all specifications: importers hold less bargaining power when their supplier has a stronger outside option. Specifically, an increase in the supplier’s past sales to other buyers relative to the importer’s purchases from other suppliers is consistently associated with lower bargaining power for the importer.
@@
We assess the robustness of the structural estimates to alternative sample definitions and model calibrations. Table 5 reports the results.
Table 5: Robustness of Model Estimates
Baseline Only Consumption All Products $\rho=5$ $\varrho=0.5$ (1) (2) (3) (4) (5) Bargaining power ($\phi$) 0.827 0.718 0.765 0.877 0.856 (0.007) (0.005) (0.004) (0.004) (0.007) Returns to scale ($\theta$) 0.455 0.558 0.508 0.465 0.470 (0.004) (0.003) (0.002) (0.002) (0.003) Observations 3,120,000 3,024,000 6,143,000 3,120,000 3,120,000
Notes: This table reports robustness checks for the structural estimates. Columns vary either the sample definition or the calibrated parameters. “Only Consumption” restricts the sample to consumption goods. “All Products” includes all product categories. Columns $\rho=5$ and $\varrho=0.5$ modify the calibrated elasticities. All columns report GMM estimates without fixed effects. All estimations use CIF duty-exclusive unit values. Observation counts are rounded in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Restricting the sample to consumption goods only (column (2)) yields a slightly lower estimate of $\phi$ and a slightly higher estimate of $\theta$. Column (3) reports estimates for the pooled sample, including all products, which lie in between, as expected. The model’s core mechanisms, therefore, operate in consumption goods transactions as well. We nonetheless retain the baseline focus on capital and intermediate inputs, for which the framework is more naturally suited. [^27] Adapting the framework specifically to consumption goods would likely require modifications to the modeling of the importer’s technology and its calibration.
On the calibration side, we consider two alternative parameter values. First, we lower the elasticity of substitution across foreign varieties to $\rho=5$ (instead of 10), at the lower end of the range reported in Anderson2004. Second, we introduce decreasing returns to scale in downstream production by setting $\varrho=0.5$, close to the supplier-side estimate. Both $\rho$ and $\varrho$ affect how cross-sectional price dispersion varies with the exporter’s supplier share, and thus with oligopoly power. Columns (4) and (5) show that the structural estimates remain highly stable. [^28] Results are robust to alternative related-party trade classifications. See Appendix D.3 and Table D.6 for details. This suggests that $\phi$ is primarily identified from variation in buyer share rather than supplier share.
@@
Using equation (2.10), we compute markups for all buyer-supplier–product matches given the estimated parameters and the observed market shares.
The resulting markup distribution is clustered near the competitive benchmark, with a median of 1.01. This contrasts with the oligopoly benchmark, whose median is 1.21 and whose interquartile range spans 1.11 to 2.45. In the model, observed markups are a convex combination of oligopoly and oligopsony markups, where the median oligopsony markup is 0.95. The large estimated bargaining power of buyers ($\hat{\phi}\approx 0.8$) shifts weight toward oligopsony, allowing importers to capture a substantial share of the surplus from exporters.